# Partial Semigroupoid Actions on Sets

### v1

@@ 2. Global and Partial Actions
Fix a functor $F\colon\mathcal{C}\to\mathcal{D}$ and $X\in\mathcal{D}_{0}$. We define a category $(X\to F)$ in the following way:
  • $(X\to F)_{0}$ is the family of morphisms $f\in\mathcal{D}_{1}$ such that $D(f)=X$ and $R(f)\in F_{0}(\mathcal{C}_{0})$. That is, the objects of $(X\to F)$ are the morphisms $f\colon X\to F_{0}(Y)$ for some $Y\in\mathcal{C}_{0}$.
  • Given $f\colon X\to F_{0}(Y)$ and $g\colon X\to F_{0}(Z)$ in $(X\to F)_{0}$, define $(X\to F)(f,g)$ as the family of morphisms $\varphi\in\mathcal{D}_{1}(F_{0}(Y)),F_{0}(Y^{\prime}))$ such that $\varphi\circ f=g$. Now, the family of morphisms $(X\to F)_{1}$ is the union $\bigcup_{f,g\in(X\to F)_{0}}(X\to F)(f,g)$.
  • The composition in $(X\to F)$ is defined as the composition in $\mathcal{D}_{1}$, according to the diagram:
    $X$$F_{0}(Y_{1})$$F_{0}(Y_{2})$$F_{0}(Y_{3})$$f$$g$$h$$\varphi$$\psi$$\psi\circ\varphi$$\varphi\circ f=g,$$\psi\circ g=h,$$(\psi\circ\varphi)\circ f=\psi\circ g=h.$
Then $(X\to F)=((X\to F)_{0},(X\to F)_{1},D,R,\circ)$ is a category, where $D$ and $R$ are domain and range functions of the category $\mathcal{D}$. In particular, if we fix $(\alpha,X)\in\mathcal{A}_{p}(S)$ a partial action and $I\colon\mathcal{A}(S)\to\mathcal{A}_{p}(S)$ the inclusion functor, given by $I_{0}(\beta,Y)=(\beta,Y)$ for every global action $(\beta,Y)$ of $S$, and $I_{1}(f)=f$ for every morphism between global actions, then $((\alpha,X)\to I)$ is a category.
@@
    If $(\beta,Y,\delta)$ is a universal globalization for $(\alpha,X)$, then $\delta$ is an initial object in $((\alpha,X)\to I)$.
@@
    Suppose that $(\alpha,X)$ admits a universal globalization. If $\gamma\colon(\alpha,X)\to(\theta,Z)$ is an initial object in $((\alpha,X)\to I)$, then $(\theta,Z,\gamma)$ is a universal globalization for $(\alpha,X)$.
Therefore, if $(\alpha,X)$ has a universal globalization, then the family of universal globalizations of $(\alpha,X)$ is precisely the family of initial objects in the category $((\alpha,X)\to I)$.
@@
(i) This is straightforward from the definition. An initial object in $((\alpha,X)\to I)$ is a morphism of partial actions $\delta\colon(\alpha,X)\to I(\beta,Y)$ such that, for each morphism $\varphi\colon(\alpha,X)\to I(\theta,Z)$, there is a unique morphism $\Phi\colon(\beta,Y)\to(\theta,Z)$ such that $\Phi\circ\delta=\gamma$. Hence, if $(\beta,Y,\delta)$ is a universal globalization, then $\delta$ is an initial object in $((\alpha,X)\to I)$.
(ii) If $\gamma\colon(\alpha,X)\to I(\theta,Z)$ is an initial object in $((\alpha,X)\to I)$, then for each morphism of partial actions $\varphi\colon(\alpha,X)\to I(\theta^{\prime},Z^{\prime})$, there is a unique morphism $\Phi\colon(\theta,Z)\to(\theta^{\prime},Z^{\prime})$ such that $\Phi\circ\gamma=\gamma^{\prime}$. Therefore, for $(\theta,Z,\gamma)$ to be a universal globalization, it remains to show that $(\theta,Z,\gamma)$ is, in fact, a globalization for $(\alpha,X)$.
For it, suppose that $(\beta,Y,\delta)$ is a universal globalization of $(\alpha,X)$. From (i) we know that $\delta$ is an initial object in $((\alpha,X)\to I)$. Since initial objects are unique up to isomorphism, there is an isomorphism $\Psi\colon\delta\to\gamma$. That is, $\Psi\colon(\beta,Y)\to(\theta,Z)$ is a bijective embedding such that $\Psi\circ\delta=\gamma$. Since $(\beta,Y,\delta)$ is a globalization, the morphism $\delta$ is an embedding. Since the composition of embeddings is again an embedding, it follows that $\Psi\circ\delta=\gamma\colon(\alpha,X)\to(\theta,Z)$ is an embedding. Therefore, $(\theta,Z,\gamma)$ is a globalization of $(\alpha,X)$ and, hence, a universal globalization. ∎
In the next section, we will prove that any partial action of $S$ admits universal globalization. Then, we obtain from Proposition 2.23 that universal globalizations of a partial action $(\alpha,X)$ are, in fact, the initial objects of the category $((\alpha,X)\to I)$. This is a generalization of [17, Corollary 3.6] for semigroups and the first part of [16, Theorem 3.4] for groups. Notice that, in the latter, it is stated that a universal globalization is an initial object in the category Global, which is the full subcategory of $((\alpha,X)\to I)$ whose objects are the embeddings $(\alpha,X)\to I(\beta,Y)$.
@@ 3. The universal globalization
Let $s,t\in S$. Then $s\sim t$ if and only if one of the following holds:
@@
The relation $\sim$ is reflexive and transitive. Reflexivity follows from (R1). To see that $s\sim t$ and $t\sim r$ imply $s\sim r$, it is enough to consider that the relations are given by (R2) or (R3). If both relations are given by (R2), then $S^{s}=S^{t}=S^{r}\neq\emptyset$. Hence $s\sim r$ by (R2). If $s\sim t$ is given by (R2) and $t\sim r$ is given by (R3), then $t=ur$ for some $u\in S$. Therefore, $S^{r}=S^{ur}=S^{t}=S^{s}\neq\emptyset$, so $s\sim r$ by (R2). Analogously, if $s\sim t$ is given by (R3) and $t\sim r$ is given by (R2), then $s=ut$ for some $u\in S$ and $S^{s}=S^{ut}=S^{t}=S^{r}\neq\emptyset$. Finally, if both relations are given by (R3), then there exist $u,v\in S$ such that $s=ut$ and $t=vr$, concluding that $s=(uv)r$ and $s\sim r$ by (R3).
The relation $\sim$ defined above may not be symmetric. For example, consider the semigroupoid $S=\{s,t,st\}$ given in Example 2.7. Then, $st\sim t$ by (R3). However, $t\not\sim st$ because $t\neq st$, $S^{t}=\emptyset$, and ${{}^{st}S}=\emptyset$.
@@
(i) Let $(s,t)\in S^{(2)}$. Then $s\in{{}^{t}S}$ and $st=s\cdot t$, so condition (R3) implies $stRt$.
(iii) We proceed by induction on the length of the chain connecting $s$ to $t$. If $s=t$, the result is trivial. If $s\neq t$ and $n=2$, then $sRt$ means that $s\sim t$ or $t\sim s$. If the relation is given by (R2), then (G1) implies that ${{}_{s}X}={{}_{t}X}$. If the relation is given by (R3), then there is $u\in S$ such that $s=ut$ or $t=us$. In any case, (G2) implies that ${{}_{s}X}={{}_{ut}X}={{}_{t}X}$ or ${{}_{t}X}={{}_{us}X}={{}_{s}X}$.
Now, assume that $s$ and $t$ are connected by a chain of length $n>2$ and suppose that, whenever $u,v\in S$ are connected by a chain of length $m<n$, then ${{}_{u}X}={{}_{v}X}$. Let $r_{1},\dots,r_{n}\in S$ be the chain connecting $s$ to $t$. Then, by the proof of $n=2$, we have ${{}_{s}X}={{}_{r_{2}}X}$. Since $r_{2},\dots,r_{n}\in S$ is a chain of length $n-1$ connecting $r_{2}$ to $t$, it follows from the induction hypothesis that ${{}_{r_{2}}X}={{}_{t}X}$. Therefore, ${{}_{s}X}={{}_{r_{2}}X}={{}_{t}X}$.
(ii) Is a consequence of (iii) and Example 2.12. ∎
@@
Consider $\approx$ as the smallest equivalence relation on $D$ that contains the relations:
@@
    $(s,x)\sim(t,y)$, if there is $u\in S^{t}$ such that $s=tu$, $x\in{{}_{u}X}$, and $\alpha_{u}(x)=y$.
More precisely, $(a,x)\approx(b,y)$ if and only if $(a,x)=(b,y)$ or there exist $(c_{1},z_{1}),\dots(c_{n},z_{n})\in D$ such that $(a,x)=(c_{1},z_{1})$, $(c_{n},z_{n})=(b,y)$ and, for each $i=1,\dots,n-1$, either $(c_{i},z_{i})\sim(c_{i+1},z_{i+1})$ or $(c_{i+1},z_{i+1})\sim(c_{i},z_{i})$. In this case, we say that there is a chain of length $n$ connecting $(a,x)$ to $(b,y)$.
@@
If $n=2$, then $(s,x)\approx(t,y)$ is given by condition (2). Suppose there exists $u\in S^{t}$ such that $s=tu$, $x\in{{}_{u}X}$, and $\alpha_{u}(x)=y$. If $x\in{{}_{s}X}={{}_{tu}X}$ then. from (P1), we obtain that $x\in{{}_{tu}X}\cap{{}_{u}X}=\alpha_{u}^{-1}(_{t}X\cap X_{u})$. Hence, $y=\alpha_{u}(x)\in{{}_{t}X}$. On the other hand, if $y\in{{}_{t}X}$, then again from (P1), we have $\alpha_{u}(x)\in{{}_{t}X}\cap X_{u}$, which implies that $x\in\alpha_{u}^{-1}(_{t}X\cap X_{u})\subseteq{{}_{s}X}$. In both cases, $\alpha_{t}(y)=(\alpha_{t}\circ\alpha_{u})(x)=\alpha_{s}(x)$. The case $t=su$ and $\alpha_{u}(y)=x$ is analogous.
If $n=3$, then $(s,x)\approx(t,y)$ is given either by (2,2) as $(s,x)\sim(r,z)\sim(t,y)$, where $r\in S$, or by (1,1) as $(s,x)\sim(\delta,z)\sim(t,y)$. The first case is obtained applying the argument for $n=2$ twice. In the second case the result is immediate, since $(s,x)\sim(\delta,z)$ and $(\delta,z)\sim(t,y)$ imply that $x\in{{}_{s}X}$, $y\in{{}_{t}X}$, and $\alpha_{s}(x)=z=\alpha_{t}(y)$.
Now, assume $n>3$ and that (ii) is valid whenever the chain connecting $(s,x)$ to $(t,y)$ has length $m<n$. If the chain starts with relation (2), we apply the case $n=2$ for the first relation and use the induction hypothesis on the remaining chain, which starts with an element $(r,z)$ with $r\in S$ and has length $n-1$. If the chain starts with (1,1), apply the argument on $n=3$ for the first two relation and the induction hypothesis for the remaining chain, of length $n-2$. If the chain starts with (1,=) it can be shortened and, hence, the result follows from the induction hypothesis.
@@
If $n=2$, then $(s,x)\sim(t,y)$ is given by (2). If $s=tu$, then $ps=(pt)u$, therefore, condition (2) implies $(ps,x)\sim(pt,y)$. The case $t=su$ is analogous. If $n=3$ then $(s,x)\approx(t,y)$ is given by (2,2) or (1,1). In the first case we have $(s,x)\sim(r,z)\sim(t,y)$ for some $r\in S$. Since both $s=ru$ and $r=sv$ implies ${{}^{s}S}={{}^{r}S}$, we have $(p,s)\in S^{(2)}$ if and only if $(p,r)\in S^{(2)}$. Thus, the result follows applying the case $n=2$ twice. For the second case, note that $(p,s)\in S^{(2)}$ and $x\in{{}_{s}X}$ implies $\alpha_{s}(x)\in{{}_{\rho_{p}}X}$, hence, $(p,\alpha_{s}(x))\in D$ and $(ps,x)\sim(p,\alpha_{s}(x))$ by (2). Now, if $(s,x)\sim(\delta,z)\sim(t,y)$, is must be $\alpha_{s}(x)=z=\alpha_{t}(y)$. Therefore
@@
Now, let $n>3$ and assume that (iii) is valid whenever the chain connecting $(s,x)$ to $(t,y)$ has length $m<n$. If the chain starts with relation (2), then $c_{2}\in S$ and the argument for (2,2) shows that $(p,c_{2})\in S^{(2)}$, thus, we apply the case $n=2$ for $(s,x)\sim(c_{2},z_{2})$ and the induction hypothesis for the remaining chain. If the chain starts with relation (1), then $x\in{{}_{s}X}$ and, from (ii), we obtain $y\in{{}_{t}X}$ and $\alpha_{s}(x)=\alpha_{t}(y)$. Therefore $(ps,x)\sim(p,\alpha_{s}(x))=(p,\alpha_{t}(y))\sim(pt,y)$ as in $n=3$.
@@
We will denote by $E$ the quotient set $D/\approx$, and the equivalence class of $(a,x)\in D$ by $[a,x]\in E$. For each $s\in S$, let
@@
and denote by ${{}_{s}E}$ the quotient set ${{}_{s}D}/\approx$. Furthermore, for each $[a,x]\in{{}_{s}E}$, consider
$\beta_{s}([a,x])=\begin{cases}[st,y],&\text{if }(a,x)\approx(t,y)\in{{}_{s}D}.\\ [s,x],&\text{if }a=\delta.\end{cases}$
It follows from Lemma 3.4(i), (iii) and (iv) that $\beta_{s}\colon{{}_{s}E}\to E$ is a well-defined function.
The next example illustrates the difference between sets $S^{\delta}\times X$, $D$, and $E$. We present the sets ${{}_{s}E}$ and show that $E$ may not be equal to the union of the sets ${{}_{s}E}$ and $\beta_{s}({{}_{s}E})$. That is, viewed as a partial action, the pair $(\beta,E)$ can be degenerate.
@@
Furthermore, for $X=\{1,2,3,4\}$ and the partial action $(\alpha,X)$ defined in Example 2.13, we have:
@@
${{}_{s_{1}}E}$ =$=\{[\delta,1],[\delta,2],[\delta,3],[t_{1},1],[t_{1},3],[t_{2},2],[t_{2},3]\},$
${{}_{s_{2}}E}$ =$=\{[\delta,1],[\delta,3],[t_{2},2],[t_{2},3]\},$
${{}_{t_{1}}E}$ =$=\{[\delta,1],[\delta,2],[\delta,3]\}={{}_{t_{2}}E}={{}_{0}E}.$
Notice that $[0,3]=[s_{1}t_{1},3]=\beta_{s_{1}}([t_{1},3])\in E_{s_{1}}$. However, the class $[\delta,4]$ does not belong to any ${{}_{s}E}$ or $E_{s}$. Therefore $\cup_{s\in S}({{}_{s}E}\cup E_{s})=E\setminus\{[\delta,4]\}$.
@@
Define $\delta\colon X\to E$ by $\delta(x)=[\delta,x]$. Then:
@@
    The triple $(\delta,\beta,E)$ is a universal globalization for $(\alpha,X)$.

### v2

@@ 2. Global and Partial Actions
In this work, all functors are assumed to be covariant, therefore we will omit the term and simply refer to them as functors. Now, fix a functor $F\colon\mathcal{C}\to\mathcal{D}$ and $X\in\mathcal{D}_{0}$. We define a category $(X\downarrow F)$ in the following way:
  • $(X\downarrow F)_{0}$ is the family of morphisms $f\in\mathcal{D}_{1}$ such that $D(f)=X$ and $R(f)\in F_{0}(\mathcal{C}_{0})$. That is, the objects of $(X\downarrow F)$ are the morphisms $f\colon X\to F_{0}(Y)$ for some $Y\in\mathcal{C}_{0}$.
  • Given $f\colon X\to F_{0}(Y)$ and $g\colon X\to F_{0}(Z)$ in $(X\downarrow F)_{0}$, define $(X\downarrow F)(f,g)$ as the family of morphisms $\varphi\in\mathcal{D}_{1}(F_{0}(Y)),F_{0}(Y^{\prime}))$ such that $\varphi\circ f=g$. Now, the family of morphisms $(X\downarrow F)_{1}$ is the union $\bigcup_{f,g\in(X\downarrow F)_{0}}(X\downarrow F)(f,g)$.
  • The composition in $(X\downarrow F)$ is defined as the composition in $\mathcal{D}_{1}$, according to the diagram:
[…]
Then $(X\downarrow F)=((X\downarrow F)_{0},(X\downarrow F)_{1},D,R,\circ)$ is a category, where $D$ and $R$ are domain and range functions of the category $\mathcal{D}$. This category is called the category of objects $F$-under $X$, and appears in [22, §6] as a particular case of a comma category. In particular, if we fix $(\alpha,X)\in\mathcal{A}_{p}(S)$ a partial action and $I\colon\mathcal{A}(S)\to\mathcal{A}_{p}(S)$ the inclusion functor, given by $I_{0}(\beta,Y)=(\beta,Y)$ for every global action $(\beta,Y)$ of $S$, and $I_{1}(f)=f$ for every morphism between global actions, then $((\alpha,X)\downarrow I)$ is a category.
@@
    If $(\iota,\beta,Y)$ is a universal globalization for $(\alpha,X)$, then $\iota$ is an initial object in $((\alpha,X)\downarrow I)$.
@@
    Suppose that $(\alpha,X)$ admits a universal globalization. If $\gamma\colon(\alpha,X)\to(\theta,Z)$ is an initial object in $((\alpha,X)\downarrow I)$, then $(\gamma,\theta,Z)$ is a universal globalization for $(\alpha,X)$.
Therefore, if $(\alpha,X)$ has a universal globalization, then the family of universal globalizations of $(\alpha,X)$ is precisely the family of initial objects in the category $((\alpha,X)\downarrow I)$.
@@
(i) This is straightforward from the definition. An initial object in $((\alpha,X)\downarrow I)$ is a morphism of partial actions $\iota\colon(\alpha,X)\to I(\beta,Y)$ such that, for each morphism $\varphi\colon(\alpha,X)\to I(\theta,Z)$, there is a unique morphism $\Phi\colon(\beta,Y)\to(\theta,Z)$ such that $\Phi\circ\iota=\gamma$. Hence, if $(\iota,\beta,Y)$ is a universal globalization, then $\iota$ is an initial object in $((\alpha,X)\downarrow I)$.
(ii) If $\gamma\colon(\alpha,X)\to I(\theta,Z)$ is an initial object in $((\alpha,X)\downarrow I)$, then for each morphism of partial actions $\varphi\colon(\alpha,X)\to I(\theta^{\prime},Z^{\prime})$, there is a unique morphism $\Phi\colon(\theta,Z)\to(\theta^{\prime},Z^{\prime})$ such that $\Phi\circ\gamma=\gamma^{\prime}$. Therefore, for $(\gamma,\theta,Z)$ to be a universal globalization, it remains to show that $(\gamma,\theta,Z)$ is, in fact, a globalization for $(\alpha,X)$.
To prove this, suppose that $(\iota,\beta,Y)$ is a universal globalization of $(\alpha,X)$. From (i) we know that $\iota$ is an initial object in $((\alpha,X)\downarrow I)$. Since initial objects are unique up to isomorphism, there is an isomorphism $\Psi\colon\iota\to\gamma$. That is, $\Psi\colon(\beta,Y)\to(\theta,Z)$ is a bijective embedding such that $\Psi\circ\iota=\gamma$. Since $(\iota,\beta,Y)$ is a globalization, the morphism $\iota$ is an embedding. Since the composition of embeddings is again an embedding, it follows that $\Psi\circ\iota=\gamma\colon(\alpha,X)\to(\theta,Z)$ is an embedding. Therefore, $(\gamma,\theta,Z)$ is a globalization of $(\alpha,X)$ and, hence, a universal globalization. ∎
Remark 2.24.
In the next section, we will prove that any partial action of $S$ admits universal globalization. Then, we obtain from Proposition 2.23 that universal globalizations of a partial action $(\alpha,X)$ are, in fact, the initial objects of the category $((\alpha,X)\downarrow I)$. This is a generalization of [19, Corollary 3.6] for semigroups and the first part of [17, Theorem 3.4] for groups. Notice that, in the latter, it is stated that a universal globalization is an initial object in the category Global, which is the full subcategory of $((\alpha,X)\downarrow I)$ whose objects are the embeddings $(\alpha,X)\to I(\beta,Y)$.
@@ 3. The universal globalization
Let $s,t\in S$. We write $s\sim t$ if and only if one of the following holds:
@@
The relation $\sim$ is reflexive and transitive. Reflexivity follows from (R1). To see that $s\sim t$ and $t\sim r$ imply $s\sim r$, it is enough to consider that the relations are given by (R2) or (R3). If both relations are given by (R2), then $S^{s}=S^{t}=S^{r}\neq\emptyset$. Hence $s\sim r$ by (R2). If $s\sim t$ is given by (R2) and $t\sim r$ is given by (R3), then $t=ur$ for some $u\in S$. Therefore, $S^{r}=S^{ur}=S^{t}=S^{s}\neq\emptyset$, and hence $s\sim r$ by (R2). Analogously, if $s\sim t$ is given by (R3) and $t\sim r$ is given by (R2), then $s=ut$ for some $u\in S$ and $S^{s}=S^{ut}=S^{t}=S^{r}\neq\emptyset$. Finally, if both relations are given by (R3), then there exist $u,v\in S$ such that $s=ut$ and $t=vr$, concluding that $s=(uv)r$ and $s\sim r$ by (R3).
The relation $\sim$ defined above may not be symmetric. For example, consider the semigroupoid $S=\{s,t,st\}$ given in Example 2.7. Then, $st\sim t$ by (R3). However, $t\not\sim st$ because $t\neq st$, $S^{t}=\emptyset$, and $^{st}S$$\emptyset$.
@@
(i) Let $(s,t)\in S^{(2)}$. Then $s\in{{}^{t}S}$ and $st=s\cdot t$, hence, condition (R3) implies $stRt$.
(iii) Let $(\alpha,X)$ be any global action of $S$ and suppose that $sRt$. We proceed by induction on the length of the chain connecting $s$ to $t$. If $s=t$, the result is straightforward. If $s\neq t$ and $n=2$, then $sRt$ means that $s\sim t$ or $t\sim s$. If the relation is given by (R2), then (G1) implies that ${{}_{s}X}={{}_{t}X}$. If the relation is given by (R3), then there is $u\in S$ such that $s=ut$ or $t=us$. In any case, (G2) implies that ${{}_{s}X}={{}_{ut}X}={{}_{t}X}$ or ${{}_{t}X}={{}_{us}X}={{}_{s}X}$.
Now, assume that $s$ and $t$ are connected by a chain of length $n>2$ and suppose that, whenever $u,v\in S$ are connected by a chain of length $m<n$, then ${{}_{u}X}={{}_{v}X}$. Let $r_{1},\dots,r_{n}\in S$ be the chain connecting $s$ to $t$. Then, by the proof of $n=2$, we have ${{}_{s}X}={{}_{r_{2}}X}$. Since $r_{2},\dots,r_{n}\in S$ is a chain of length n1n-1 connecting $r_{2}$ to $t$, it follows from the induction hypothesis that ${{}_{r_{2}}X}={{}_{t}X}$. Therefore, ${{}_{s}X}={{}_{r_{2}}X}={{}_{t}X}$.
(ii) Follows from (iii) applied to the global action provided in Example 2.12. ∎
@@
Consider $\approx$ to be the smallest equivalence relation on $D$ that contains the relations:
@@
    $(s,x)\sim(t,y)$, if there exists $u\in S^{t}$ such that $s=tu$, $x\in{{}_{u}X}$, and $\alpha_{u}(x)=y$.
More precisely, $(a,x)\approx(b,y)$ if and only if $(a,x)=(b,y)$, or there exist elements $(c_{1},z_{1}),\dots(c_{n},z_{n})\in D$ such that $(a,x)=(c_{1},z_{1})$, $(c_{n},z_{n})=(b,y)$ and, for each $i=1,\dots,n-1$, either $(c_{i},z_{i})\sim(c_{i+1},z_{i+1})$ or $(c_{i+1},z_{i+1})\sim(c_{i},z_{i})$. In this case, we say that there is a chain of length $n$ connecting $(a,x)$ to $(b,y)$.
@@
If $n=2$, the relation $(s,x)\sim(t,y)$ is given by condition (2). Therefore $s=tu$ or $t=su$, for some $u\in S$. In the first case it must be $x\in{{}_{u}X}$ and $\alpha_{u}(x)=y$. Thus, if $x\in{{}_{s}X}={{}_{tu}X}$, then (P1) implies that $x\in{{}_{tu}X}\cap{{}_{u}X}=\alpha_{u}^{-1}(_{t}X\cap X_{u})$. Hence, $y=\alpha_{u}(x)\in{{}_{t}X}$. On the other hand, if $y\in{{}_{t}X}$, then again from (P1), we have $\alpha_{u}(x)\in{{}_{t}X}\cap X_{u}$, which implies that $x\in\alpha_{u}^{-1}(_{t}X\cap X_{u})\subseteq{{}_{s}X}$. In both cases, $\alpha_{t}(y)=(\alpha_{t}\circ\alpha_{u})(x)=\alpha_{s}(x)$. The case $t=su$ and $\alpha_{u}(y)=x$ is analogous.
If $n=3$, then $(s,x)\sim(a,z)\sim(t,y)$, where either $a\in S$ and both relations are given by (2), or $a=\delta$ and both relations are given by (1). In the first case, we can apply the argument for $n=2$ twice. In the second case, the claim follows directly, since $(s,x)\sim(\delta,z)$ and $(\delta,z)\sim(t,y)$ imply that $x\in{{}_{s}X}$, $y\in{{}_{t}X}$, and $\alpha_{s}(x)=z=\alpha_{t}(y)$.
Now, assume $n>3$ and that (ii) holds whenever the chain connecting $(s,x)$ to $(t,y)$ has length $m<n$. Since $n>3$, we can write the chain as
$(s,x)\sim(a,z)\sim(b,w)\sim\dots\sim(t,y).$
If the relation $(s,x)\sim(a,z)$ is given by (2), then $a\in S$. By the case $n=2$, we have that $x\in{{}_{s}X}$ if and only if $z\in{{}_{a}X}$, and in this case $\alpha_{s}(x)=\alpha_{a}(z)$. Furthermore, since $a\in S$ and the chain connecting $(a,z)$ to $(t,y)$ has length n1n-1, the induction hypothesis gives that $z\in{{}_{a}X}$ if and only if $y\in{{}_{t}X}$, and in this case $\alpha_{a}(z)=\alpha_{t}(y)$. Therefore, $x\in{{}_{s}X}$ if and only if $z\in{{}_{a}X}$ if and only if $y\in{{}_{t}X}$, and in this case $\alpha_{s}(x)=\alpha_{a}(w)=\alpha_{t}(y)$.
If, on the other hand, the relation $(s,x)\sim(a,z)$ is given by (1), then $a=\delta$ and the relation $(a,z)\sim(b,w)$ is also given by (1). Hence $b\in S$. By the case $n=3$, we obtain that $x\in{{}_{s}X}$ if and only if $w\in{{}_{b}X}$, and in this case $\alpha_{s}(x)=\alpha_{b}(w)$. Since $b\in S$ and the chain connecting $(b,w)$ to $(t,y)$ has length n2n-2, the induction hypothesis yields that $w\in{{}_{b}X}$ if and only if $y\in{{}_{t}X}$, and in this case $\alpha_{b}(w)=\alpha_{t}(y)$. Therefore, $x\in{{}_{s}X}$ if and only if $w\in{{}_{b}X}$ if and only if $y\in{{}_{t}X}$, and in this case $\alpha_{s}(x)=\alpha_{b}(w)=\alpha_{t}(y)$. This completes the proof of (ii).
@@
If $n=2$, then $(s,x)\sim(t,y)$ is given by (2). If $s=tu$, then $ps=(pt)u$, therefore, condition (2) implies $(ps,x)\sim(pt,y)$. The case $t=su$ is analogous.
If $n=3$, then $(s,x)\sim(a,z)\sim(t,y)$ where either $a\in S$ and both relation are given by (2) or $a=\delta$ and both relations are given by (1). In the first case we have $s=au$ or $a=sv$, both which implies ${{}^{s}S}={{}^{a}S}$. Thus, the result follows applying the case $n=2$ twice. For the second case, note that $(p,s)\in S^{(2)}$ and $x\in{{}_{s}X}$ implies $\alpha_{s}(x)\in{{}_{\rho_{p}}X}$, hence, $(p,\alpha_{s}(x))\in D$ and $(ps,x)\sim(p,\alpha_{s}(x))$ by (2). On the other hand, if $(s,x)\sim(\delta,z)\sim(t,y)$, then is must be $\alpha_{s}(x)=z=\alpha_{t}(y)$. Therefore,
@@
Now, let $n>3$ and assume that (iii) is valid whenever the chain connecting $(s,x)$ to $(t,y)$ has length $m<n$. Since $n>3$, we can write the chain as
$(s,x)\sim(a,z)\sim(b,w)\sim\dots\sim(t,y).$
If the chain starts with relation (2), then $a\in S$, and the case $n=2$ shows that $(p,a)\in S^{(2)}$. Hence, by the result for $n=2$, we have $(ps,x)\approx(pa,z)$. Since $a\in S$, $(p,a)\in S^{(2)}$ and the chain connecting $(a,z)$ to $(t,y)$ has length n1n-1, the induction hypothesis gives $(pa,z)\approx(pt,y)$. Therefore, $(ps,x)\approx(pa,z)\approx(pt,y)$.
If, on the other hand, the chain starts with relation (1), then $x\in{{}_{s}X}$. By (ii), we obtain $y\in{{}_{t}X}$ and $\alpha_{s}(x)=\alpha_{t}(y)$. Hence $(ps,x)\sim(p,\alpha_{s}(x))=(p,\alpha_{t}(y))\sim(pt,y)$, as in the case $n=3$.
@@
We will denote by $E$ the quotient set $D/\approx$, and the equivalence class of $(a,x)\in D$ by $a,x]\in E$. For each $s\in S$, let
@@
and denote by $_{s}E$} the quotient set $_{s}D$$\approx$. Furthermore, for each $a,x]\in{{}_{s}E}$, define
$\beta_{s}([a,x])=\begin{cases}[st,y],&\text{if }(a,x)\approx(t,y)\in{{}_{s}D},\\ [s,x],&\text{if }a=\delta.\end{cases}$
We now show that $\beta_{s}\colon{{}_{s}E}\to E$ is a well-defined function. Let $a,x]\in{{}_{s}E}$. First, we prove that $\beta_{s}([a,x])\in E$. Indeed, if $(a,x)\approx(t,y)$ and $st$ is defined, then $(st,y)\in D$ by Lemma 3.4(i). On the other hand, if $(\delta,x)\in{{}_{s}D}$, then $x\in{{}_{\rho_{s}}X}$. Hence, $(s,x)\in D$ by definition. In both cases, we conclude that $\beta_{s}([a,x])\in{D/\approx}=E$. Next, we show that $\beta_{s}([a,x])$ does not depend on the choice of the representatives. Suppose that $(a,x)\approx(t,y)$ and $(a,x)\approx(r,z)$, with both $st$ and $sr$ defined. Then $[st,y]=[sr,z]$ by Lemma 3.4(iii), which implies that $\beta_{s}([a,x])$ is independent of the representative $(t,y)$. Finally, assume that $a=\delta$ and $(\delta,x)\approx(t,y)$ for some $(t,y)\in D$ with $st$ defined. By Lemma 3.4(iv), we have $[s,x]=[st,y]$, showing that both definitions of $\beta_{s}([a,x])$ coincide whenever ambiguity could arise. Hence, $\beta_{s}\colon{{}_{s}E}\to E$ is well defined.
The next example illustrates the difference between the sets $S^{\delta}\times X$, $D$, and $E$. In particular, we show that $E$ may not be equal to the union of the sets $_{s}E$} and $\beta_{s}({{}_{s}E})$. That is, viewed as a partial action, the pair $(\beta,E)$ can be degenerate.
@@
For $X=\{1,2,3,4\}$ and the partial action $(\alpha,X)$ defined in Example 2.13, we have:
@@
${{}_{s_{1}}E}$ $=\{[\delta,1],[\delta,2],[\delta,3],[t_{1},1],[t_{1},3],[t_{2},2],[t_{2},3]\},$
${{}_{s_{2}}E}$ $=\{[\delta,1],[\delta,3],[t_{2},2],[t_{2},3]\},$
${{}_{t_{1}}E}$ $=\{[\delta,1],[\delta,2],[\delta,3]\}={{}_{t_{2}}E}={{}_{0}E}.$
Notice that $0,3]=[s_{1}t_{1},3]=\beta_{s_{1}}([t_{1},3])\in E_{s_{1}}$. However, the class $\delta,4]$ does not belong to any $_{s}E$} or $E_{s}$. Therefore $\cup_{s\in S}({{}_{s}E}\cup E_{s})=E\setminus\{[\delta,4]\}$.
@@
Define $\iota\colon X\to E$ by $\iota(x)=[\delta,x]$. Then:
@@
    The triple $(\iota,\beta,E)$ is a universal globalization for $(\alpha,X)$.